A boy has twice as much money as his sister. If he gives P10 to his sister, they would have the sake amount. How much money did each have?
Answer by ikleyn(52817)

.
A boy has twice as much money as his sister. If he gives P10 to his sister, they would have the sake amount.
How much money did each have?
~~~~~~~~~~~~


x = sister.

2x = boy.


After giving P10 to the sister


    2x - 10 = x + 10


Simplify and solve


    2x - x = 10 + 10

       x   =    20


ANSWER.  Sister has P20;  boy has twice as much, i.e.  P40.


CHECK.   After giving  40 - 10 = 30 = 20 + 10.    ! Correct !

Solved.